Suppose you travel 66 kilometers in 1.25 hours. Moving at the same speed, how many kilometers would you cover in 2 hours?
Answer by checkley77(12569)

1.25/66=2/x
1.25x=66*2
1.25x=132
x=132/1.25
x=105.6 miles.